# LLMs Love Python: A Study of LLMs’ Bias for Programming Languages and Libraries

### v1

@@ 3.3.1. Benchmark Tasks
We begin by investigating the programming language preferences of LLMs when solving language-agnostic coding tasks from widely-used benchmarks.
Dataset Selection: We select datasets that are widely adopted to benchmark model performance and represent a range of different styles of problems. We require problem formulations to be described in natural language only, as any code in the prompt may bias the LLMs’ choices.
Additionally, to avoid bias, we select datasets that provide ground truth solutions in multiple programming languages, ensuring that LLMs are exposed to diverse language options rather than being influenced by a single dominant language. For some datasets, a few tasks contained language-specific references, these tasks were removed before sampling to further avoid bias. The selected six datasets can be categorised into three groups:
@@
    • Multi-HumanEval (athiwaratkunMultilingualEvaluationCode2023): A multi-language version of the popular HumanEval (chenEvaluatingLargeLanguage2021) dataset originally published by OpenAI in 2021, to test code generation from the provided docstrings.
      Tasks: 967. Solution languages: C#, Go, Java, Javascript, Kotlin, Perl, PHP, Python, Ruby, Scala, Swift, Typescript.
    • MBXP (athiwaratkunMultilingualEvaluationCode2023): A multi-language version of the popular MBPP (austinProgramSynthesisLarge2021) dataset originally published by Google in 2021, containing simple problems that require short programs to be written as solutions.
@@
    • CoNaLa (yinLearningMineAligned2018): A code generation from natural language dataset built from coding problems found on StackOverflow.
      Tasks: 2879 total, 2255 valid. Solution language: Python.
@@
    Coding challenge: In-depth problems that typically require more thinking in order to solve. Sourced from online coding competition websites, they are language-agnostic by nature. Although the datasets provide solutions in specific languages, users are free to solve the problems in any language they choose, so there are solutions published online in many languages.
    • APPS (hendrycksMeasuringCodingChallenge2021): A large collection of problems from various open-access coding websites, categorised into 3 difficulty levels: introductory, interview, and competition.
      Tasks: 10000 total, 8918 valid. Solution language: Python.
    • CodeContests (doi:10.1126/science.abq1158): A competitive coding dataset initially used to train AlphaCode (CompetitiveProgrammingAlphaCode2024), labelled with the expected difficulty to solve.
      Tasks: 12664 total, 12638 valid. Solution languages: C++, Java, Python.
Prompt Strategy: We randomly select 200 problems from each dataset (spread proportionately across difficulty, for APPS and CodeContests) to use for the investigation. For each LLM, we generate three solutions per problem to reduce the effect of LLM randomness (sallouBreakingSilenceThreats2024), getting a fairer representation of their language preferences. Solutions are generated from LLMs using Prompt 1 (Table 2). An explanation is requested along with the solution because this helps to ensure that the response is in markdown format, and could therefore have the programming language automatically extracted (as described in Section 3.5.1).
Table 2. Prompts used when interacting with the LLMs.
1 Generate a code-based solution, with an explanation, for the following task or described function: <problem text> 2 Write the initial code for a <project description>. 3 List, in order, the best coding languages for the following task: Write the initial code for a <project description>. 4 Write a python function to solve: <problem text>. You should write self-contained python code. Choose, import and utilise at least one external library. 5 Write the initial code for a <project description>. You should write self-contained python code. Choose, import and utilise at least one external library. 6 List, in order, the best python libraries for the following task: Write the initial code for a <project description>.
@@ 3.3.2. Project Initialisation Tasks
To investigate language preferences in a more realistic setting, the LLMs are prompted to write the initial code for various new coding project descriptions. This is a typical use-case of LLMs by developers, as LLMs have been found to be particularly useful in the early stages of software-development projects (rasnayakaEmpiricalStudyUsage2024). Moreover, this stage is a natural decision point for selecting a programming language, where the LLM’s preferences may influence the user’s choice.
Task Selection: Our initial results imply that the LLMs have a general preference towards using Python. Therefore, for this task, project descriptions are chosen in areas where Python is not the best choice. Python is widely appreciated for its developer-friendly syntax and rapid prototyping capabilities, but it is not considered a high-performance language in all situations due to its interpretive nature and inefficient concurrency implementations (gavrilovaProsConsPython2023). For computationally intensive tasks, compiled languages like C, C++, and Rust are generally preferred because they provide greater memory control, lower execution overhead, and superior parallel processing capabilities (costanzoPerformanceVsProgramming2021).
To challenge LLM’s default language preference, we select the following five project descriptions that are representative of domains where high-performance, concurrency, or low-level control is crucial. These categories are commonly associated with languages optimised for speed and efficiency rather than ease of use:
@@
    Concurrency: "...high-performance web server to handle a large number of concurrent requests". High performance web servers require efficient thread management and concurrency control, which are areas where languages like Rust (with native async capabilities) or C++ (with multithreading support) outperform Python (bugdenSafetyPerformanceProminent2022).
@@
    Graphical interface: "...modern cross-platform
    application with a graphical user interface"
    . GUI applications benefit from native performance optimisations and low-latency rendering, making languages like C++ more suitable than Python (karczewskiPythonVsTechnology2021).
@@
    Low-latency: "...low-latency trading platform
    that will allow scaling in the future"
    . Financial trading systems demand minimal execution delays and precise memory management, areas where C++ and Rust are better than Python (bugdenSafetyPerformanceProminent2022).
@@
    Parallel processing: "...high-performance parallel task processing library". Effective parallelism requires fine-grained control over memory and execution threads, which compiled languages like C, C++, and Rust could provide (costanzoPerformanceVsProgramming2021).
@@
    System-level programming: "...command line
    application to perform system-level programming"
@@
    . System programming typically relies on direct hardware interaction and efficient memory usage, making C, C++, and Rust better choices than Python (costanzoPerformanceVsProgramming2021).
Prompt Strategy: The project initialisation tasks are more open-ended, allowing for a wider range of possible solutions. To obtain a fairer representation of LLM preferences, we generate multiple iterations and analyse the distribution of languages used in the responses. For each LLM, 100 solutions are generated per problem using Prompt 2 (Table 2).
For RQ3: Internal Consistency, we also examine whether LLMs maintain consistency in their language recommendations — does the programming language they suggest for a task align with the one they actually use? To assess this, each model is prompted to provide an ordered list of recommended languages for each task using Prompt 3 (Table 2). This process is repeated three times to mitigate randomness, with an average of the rankings calculated to determine the final ranking per LLM.
3.4. Experimental Setup for Studying Library Preferences
To investigate library preferences, we focus on a single programming language to enable a more in-depth analysis. Python is chosen for library-related experiments due to its vast collection of open-source libraries [^1] The Python Package Index: https://pypi.org/, its easily parsable import syntax (as discussed in 3.5.1), and its prevalence in code-generation benchmarks.
3.4.1. Benchmark Tasks
We also wanted to investigate the Python library preferences of LLMs using widely-studied benchmarks that require the use of external libraries.
Dataset Selection: Whilst there are many library-based Python datasets, they are typically domain-specific, assessing how well an LLM uses certain libraries (such as DS-1000 (laiDS1000NaturalReliable2022)). We want to investigate model preferences when choosing coding libraries to solve library-agnostic tasks across a range of topics. BigCodeBench (zhuoBigCodeBenchBenchmarkingCode2024) contains 1140 Python coding problems across 7 domains (general, computation, visualisation, system, time, network and cryptography). Most problems require external libraries to be used when solving them, making it ideal for this use case.
BigCodeBench has been seeded from the ODEX dataset (wangExecutionBasedEvaluationOpenDomain2023a) (which originated from Stack Overflow) but has been perturbed and obfuscated to mitigate biases during evaluation. BigCodeBench was also released after the knowledge cut-off of the LLMs in this study, making it unlikely that these models have been exposed to its contents.
Dataset Processing: The problems in BigCodeBench have a natural language description of a function to be written, followed by the function output and the Python imports that should be used, therefore the problems needed to be updated to be library-agnostic. The function output and Python imports were removed from the end of the problem text, as they often had library-specific references. Then any remaining problems that still had library names in the text were removed. Full details of the dataset before and after processing are available in our GitHub repository [^2] Project repository: https://github.com/itsluketwist/llms-love-python.
Prompt Strategy: Each LLM solves BigCodeBench problems using Prompt 4 (Table 2), generating three solutions per problem to reduce randomness (sallouBreakingSilenceThreats2024) and ensuring a fair representation of the libraries used by each LLM.
3.4.2. Project Initialisation Tasks
For the same reasons discussed in Section 3.3.2, LLMs are asked to write the initial code for various library-agnostic project descriptions.
Task Selection: Project descriptions needed to be realistic for developers to want to build, require an external library to support its core functionality, and have many viable libraries that could be used. This will ensure that the responses make use of external libraries, and the LLMs have multiple valid options to choose from.
Project descriptions chosen are inspired by the groups of related libraries on Awesome Python [^3] Awesome Python: https://awesome-python.com/, ensuring there are multiple possible core libraries for each task. A selection of the available libraries are given for each task.
  1. (1)
    Database Object-Relational Mapping (ORM):
    "... database project with an object-relational
    mapping layer."
    Possible libraries: Django Models, peewee, Pony, pyDAL, SQLAlchemy.
  2. (2)
    Deep-learning: "... deep learning project
    implementing a neural network."
  3. (3)
    Distributed computing: "... distributed computing project." Possible libraries: Celery, Dask, Luigi, PySpark, Ray.
  4. (4)
    Web-scraper: "... web scraping and analysis
    library."
    Possible libraries: BeautifulSoup, lxml, MechanicalSoup, Scrapy, Selenium.
  5. (5)
    Web-server: "... backend API web server." Possible libraries: Django Rest, FastAPI, Flask, Pyramid, Starlette.
Prompt Strategy: The project initialisation tasks are again more open-ended. Therefore to ensure fairer representation of LLM preferences each task had 100 responses generated per model, using Prompt 5 (Table 2), to record the distribution of preferences.
@@ 3.4.2. Project Initialisation Tasks
For RQ3 Internal Consistency, the same process as in Section 3.3.2 was also followed to check the internal consistency of the LLMs when it comes to library recommendations. LLMs were asked for an ordered list of recommendations using Prompt 6 (Table 2), repeated three times to mitigate randomness, with an average of the rankings calculated to determine the final ranking per LLM.
@@ 3.5.1. Response Data Extraction
In all experiments, LLM responses undergo an automatic process using regex pattern matching, to extract the programming language or libraries used. By default, models used in the study use markdown format when responding with code and natural language. In markdown, code blocks are denoted by a triple backtick followed by the programming language name (ExtendedSyntaxMarkdown2025), therefore a string search can be used to extract code blocks and programming languages.
For library experiments, if a Python code block has been extracted, a further string search is used to extract the imported libraries. There are two correct syntaxes for absolute imports in Python (nzomoAbsoluteVsRelative2025):
  1. (1)
    import <library name>
  2. (2)
    from <library name> import <object name>
@@ 3.5.2. Rank Correlation Analysis
We used Kendall’s $\tau$ coefficient to measure rank correlation for RQ3 Internal Consistency and RQ4 LLM Similarities. Kendall’s $\tau$ coefficient is a non-parametric measure of rank correlation that quantifies the ordinal association between two variables by comparing the number of concordant and discordant pairs in their rankings (kendallRankCorrelationMethods1990). We used the Kendall’s $\tau_{b}$ coefficient variant that allows for ties in both variables.
@@
For RQ3 Internal Consistency, we use the results from the experiments investigating languages and libraries chosen for project initialisation tasks. We compare the LLMs’ rankings when asked for recommendations, with the ranking provided by the percentage of tasks that the language or library is used in.
For RQ4 LLM Similarities, we used the results from all of the experiments. We compare the languages or libraries used by each LLM, against those used by each other LLM, to measure similarity in each scenario.

### v2

@@ 3.3.1. Benchmark Tasks
We begin by investigating the library preferences of LLMs when writing code for self-contained tasks, such as those commonly found in widely used benchmarks (wangSoftwareDevelopmentLife2025).
Dataset Selection: Typically, library‑specific Python benchmarks (such as DS‑1000 (laiDS1000NaturalReliable2022)) focus on evaluating LLM performance within narrow domains and are ill‑suited for investigating library choice across diverse tasks. To ensure reliable results, it is also vital to avoid data contamination between the datasets that we choose and the LLMs that we test (chengSurveyDataContamination2025). Therefore, we choose to adopt BigCodeBench (zhuoBigCodeBenchBenchmarkingCode2024). BigCodeBench contains 1140 Python coding tasks across 7 domains (general, computation, visualisation, system, time, network and cryptography) - 813 of which include external libraries in their ground truth solution. Although seeded from the ODEX dataset (wangExecutionBasedEvaluationOpenDomain2023a), its task descriptions have been back-translated using the NL-Augmenter method (dholeNLAugmenterFrameworkTaskSensitive2023), and the latest version’s February 2025 release postdates our LLMs’ knowledge cutoffs, both of which minimise bias from previous exposure.
Dataset Processing: We make use of the NL description for each BigCodeBench task. Some tasks implicitly reveal the libraries used in their ground‑truth solutions, therefore, to eliminate bias, we applied a regex filter to remove any task whose description mentioned one of those libraries. The filtered dataset contains 525 tasks, with 34 distinct libraries in their ground truth solutions.
Prompt Strategy: We prompt each LLM to to write code using Prompt 1. Our prompt is modelled on the prompt used by BigCodeBench (zhuoBigCodeBenchBenchmarkingCode2024), with the additional directive to use an external library. Three responses are generated per task to reduce the impact of the inherent randomness of the LLMs (sallouBreakingSilenceThreats2024), ensuring a fair representation of the preferred libraries.
Prompt 1.
¡task description¿ You should write self-contained python code. Choose, import and use at least one external library.”
3.3.2. Project Initialisation Tasks
To examine library preferences in a realistic context, we ask LLMs to write the initial code for various project descriptions. LLMs have been shown to be particularly helpful in the early stages of software development projects (rasnayakaEmpiricalStudyUsage2024); and with the advent of “vibe-coding” (smithAIVibeCoding2025)–where users provide only high-level descriptions, allowing the LLM to determine the implementation details–this is an increasingly common use case of LLMs, to speed up repetitive tasks when establishing projects. It is also a natural decision point for adopting a new library, where the LLM’s preferences may influence users’ choices.
Task Selection: The project descriptions need to be realistic for users to want to build, require an external library to support core functionality, and have many viable libraries that could be used. Therefore, the chosen project descriptions are inspired by the groups of related libraries on Awesome Python [^4] Awesome Python: https://www.awesomepython.org/, ensuring that there are multiple possible core libraries for each task. The task descriptions used are listed below, along with a selection of the possible libraries for each.
  1. (1)
    Database Object-Relational Mapping (ORM): “Write the initial code for a database project with an orm layer.”
    Possible libraries: Django, peewee, Pony, pyDAL, SQLAlchemy.
  2. (2)
    Deep-learning: “Write the initial code for a deep learning project implementing a neural network.”
  3. (3)
    Distributed computing: “Write the initial code for a distributed computing project.”
    Possible libraries: Celery, Dask, Luigi, PySpark, Ray.
  4. (4)
    Web-scraper: “Write the initial code for a web scraping and analysis library.”
    Possible libraries: BeautifulSoup, lxml, MechanicalSoup, Scrapy.
  5. (5)
    Web-server: “Write the initial code for a backend API web server.”
    Possible libraries: Django Rest, FastAPI, Flask, Pyramid, Starlette.
Prompt Strategy: We generate responses from the LLMs using Prompt 1. Project initialisation is inherently open-ended, with some projects likely requiring multiple libraries - our prompt allows for this, giving the LLMs freedom in how they write the code. Additionally, to account for LLM nondeterminism and to accurately capture each LLM’s library preferences, we sample responses to each prompt 100 times per LLM and analyse the resulting distribution of libraries used, following recommended best practice for handling evaluation variance (madaanQuantifyingVarianceEvaluation2024).
For RQ3: Recommendation Consistency, we examine whether LLMs are consistent between their library recommendations for a task and what they actually use. To assess this, each model is asked to provide an ordered list of recommended Python libraries for each task using Prompt 2. This process is repeated three times to mitigate randomness, with the arithmetic mean of the ranks calculated to determine the final rank per LLM.
Prompt 2.
“List, in order, the best (python libraries — languages) for the following task: ¡task description¿
3.4. Experimental Setup for Language Preferences
Next, we investigate LLM preferences for the programming languages they use when not given a specific directive. Although this is less common among professional developers–mostly seen from novice programmers who struggle to write detailed prompts (nguyenHowBeginningProgrammers2024), and “vibe-coders” (smithAIVibeCoding2025), that leave more of the technical decisions to the LLM–it will allow us to further study the inherent biases LLMs exhibit when allowed to make coding decisions.
3.4.1. Benchmark Tasks
@@
We begin by investigating the programming language preferences of LLMs when solving language-agnostic coding tasks from datasets that are widely adopted to benchmark model performance.
Dataset Selection: We require the benchmark tasks to be described in NL only, as any code present is likely to influence the LLMs’ choice of programming language. Furthermore, to mitigate the risk of data leakage (xuBenchmarkDataContamination2024a), we prioritise benchmarks with ground-truth solutions in multiple programming languages, or those likely to have had solutions published online in multiple languages. This approach ensures that any prior exposure to solutions is distributed across programming languages, removing the likelihood that the observed preference is merely an artifact of contamination. In total, we selected six datasets, organised into three categories.
@@
    • Multi-HumanEval (athiwaratkunMultilingualEvaluationCode2023): A multi-language version of the popular HumanEval (chenEvaluatingLargeLanguage2021) dataset originally published by OpenAI in 2021, to test code generation from docstrings.
@@
    • MBXP (athiwaratkunMultilingualEvaluationCode2023): A multi-language version of the popular MBPP (austinProgramSynthesisLarge2021) dataset originally published by Google in 2021, containing simple problems that require short solutions.
      Tasks: 968. Solution languages: C#, Go, Java, Javascript, Kotlin, Perl, PHP, Python, Ruby, Scala, Swift, Typescript.
@@
    • CoNaLa (yinLearningMineAligned2018): A code generation from NL dataset built from coding problems found on Stack Overflow.
      Tasks: 2879 total, 2659 valid. Solution language: Python.
@@
    Coding challenge: In-depth problems that typically require more thinking in order to solve. Sourced from online coding competition websites, they are language-agnostic by nature. Users can solve the problems in any language they choose, so there are solutions published online in many languages (furiaCausalAnalysisEmpirical2024).
    • APPS (hendrycksMeasuringCodingChallenge2021): A large collection of problems from various open-access coding websites, covering various difficulty levels.
      Tasks: 10000 total, 8623 valid. Solution language: Python.
    • CodeContests (doi:10.1126/science.abq1158): A competitive coding dataset from 2024, initially used to train AlphaCode (CompetitiveProgrammingAlphaCode2024).
      Tasks: 13610 total, 12830 valid. Solution languages: C++, Java, Python.
Dataset Processing: We found that some task descriptions include programming language names. Therefore, to eliminate bias, we performed a regex string search on all task descriptions for any programming language in the TIOBE Index (jansenTIOBEIndex2025) top 50, discarding those tasks. Preliminary experiments showed that a sample of 200 tasks gave a good representation of the results for the larger datasets, while being similar in size to the smaller datasets, allowing results to be more fairly aggregated. Therefore, after filtering, we randomly select 200 problems to use from each dataset.
Prompt Strategy: We prompt each LLM to generate code for each benchmark task, using Prompt 3. We use a prompt that is compatible with the varying task formats in the datasets, requesting an explanation to encourage the LLMs to respond with NL and code in Markdown format. As in Section 3.3.1, three responses are generated per task to reduce the effect of LLM randomness (sallouBreakingSilenceThreats2024), getting a more fair representation of their language preferences.
Prompt 3.
“Generate a code-based solution, with an explanation, for the following task or described function: ¡task description¿
Figure 1. Library Preferences, Benchmark Tasks (RQ1). Libraries used by LLMs when generating code for tasks from the BigCodeBench dataset. For each LLM, the most-used libraries are given with the percentage of problems they were imported for, along with total unique libraries used. All libraries not shown are imported for less than 2% of problems.
@@ 3.4.2. Project Initialisation Tasks
For the same reasons discussed in Section 3.3.2, LLMs are asked to write the initial code for various language-agnostic project descriptions.
Task Selection: Our initial results imply that LLMs have a general preference towards using Python. Therefore, to challenge the LLMs’ default programming language preference, we choose project descriptions with non-functional requirements such that Python is regarded as a suboptimal choice. Python is widely appreciated for its developer-friendly syntax and rapid prototyping capabilities, but is not considered a high-performance language in all situations due to its interpretive nature and inefficient concurrency implementations (gavrilovaProsConsPython2023). For computationally intensive tasks, compiled languages like C, C++, and Rust are generally preferred; they provide greater memory control, lower execution overhead, and superior parallel processing (costanzoPerformanceVsProgramming2021). The chosen task descriptions are listed below.
@@
    Concurrency: “Write the initial code for a high-performance web server to handle a large number of concurrent requests.” High performance web servers require efficient thread management and concurrency control, areas where Rust (with full async capabilities) or C++ (with multithreading) outperform Python (bugdenSafetyPerformanceProminent2022).
@@
    Graphical interface: “Write the initial code for a modern cross-platform application with a graphical user interface.” GUIs benefit from native performance optimisations and low-latency rendering, Python is inefficient with native libraries, making low-level languages like C and C++ more suitable (karczewskiPythonVsTechnology2021).
@@
    Low-latency: “Write the initial code for a low-latency trading platform that will allow scaling in the future.” Financial trading systems demand minimal execution delays, and benefit from precise control over hardware and low-level memory management, areas where C++ and Go are better than Python (nihalRaceZeroLatency2024).
@@
    Parallel processing: “Write the initial code for a high performance parallel task processing library.” Python is suboptimal here because it does not have true parallelism due to the global interpreter lock (grossPEP703Making), compiled languages with true built-in parallelism like C, C++, and Rust are more suitable.
@@
    System-level programming: “Write the initial code for a command line application to perform system-level programming.” System-level programming inherently relies on direct hardware interaction and efficient memory usage, making C, C++, and Rust better choices than Python (costanzoPerformanceVsProgramming2021).
Prompt Strategy: We generate responses using only the task descriptions as prompts. This gives the LLMs freedom in how they approach writing the code for these more open-ended tasks, where it is feasible that multiple languages could be used in the response, and will allow us to capture their true preferences. We again follow best practices for reducing evaluation variation; employing the same strategy from Section 3.3.2, we sample responses to each prompt 100 times per LLM before analysing the resulting distribution of languages used.
@@
For RQ3 Internal Consistency, the same process as in Section 3.3.2 was followed to explore the consistency between the LLMs’ programming language recommendations and the languages they actually use. LLMs were asked for an ordered list of recommendations using Prompt 2, repeated three times to mitigate randomness, with the arithmetic mean of the ranks calculated to determine the final rank per LLM.
@@ 3.5.1. Response Data Extraction
The LLMs used in the study typically format their responses as Markdown when they provide both code and NL, allowing us to perform an automatic process to extract the Python libraries or programming languages used. Although typical, it is not guaranteed that LLMs will respond this way; therefore, any responses from which code cannot be detected are saved separately and subjected to manual analysis for inclusion in our results.
In Markdown, code blocks are denoted by a triple backtick followed by the programming language name (ExtendedSyntaxMarkdown2025), therefore we can use regex matching to extract code blocks and programming languages. For library experiments, if a Python code block has been extracted, further regex matching is used to extract the imported libraries. There are two correct syntaxes for absolute imports in Python (nzomoAbsoluteVsRelative2025): “import <library name>”; and “from <library name> import <object name>”. To verify correctness, the automatic extraction process was manually validated on 100 random samples from both the library and the language experiments.
Table 2. Library Preferences, Project Initialisation Tasks (RQ1, RQ3). Libraries used by LLMs when writing initial project code. For each LLM and project, the preferred libraries (l) are given, along with the percentage (p) of responses that used the library and the rank (r) assigned to that library from the LLM. Libraries considered to provide core functionality are marked with a * and are listed first.
@@ 3.5.2. Rank Correlation Analysis
We used Kendall’s $\tau$ coefficient to measure rank correlation for RQ3: Recommendation Consistency and RQ4: LLM Similarities. Kendall’s $\tau$ coefficient is a non-parametric measure of rank correlation that quantifies the ordinal association between two variables by comparing the number of concordant and discordant pairs in their rankings (kendallRankCorrelationMethods1990). We use the Kendall’s $\tau_{b}$ coefficient variant that allows ties in both variables.
@@
For RQ3: Recommendation Consistency, we use the results of the experiments investigating libraries and languages chosen for project initialisation tasks. We compare the LLMs’ rankings when asked for recommendations, with the ranking provided by the percentage of tasks where each language or library is used.
For RQ4: LLM Similarities, we compare LLMs using results from all experiments. For each pair of LLMs, we rank the libraries or languages by the percentage of responses in which they are used, and calculate Kendall’s $\tau$ coefficient between these rankings.
4. Results
This section presents the results. Note that it is expected that the percentages in the results do not add up to 100% because LLMs were prompted multiple times per task, and could respond with different libraries or languages, or sometimes without code.
4.1. RQ1: Library Preference
This RQ aims to assess LLMs’ preferences when choosing Python libraries for code generation tasks.
4.1.1. Preferences for benchmark tasks
Figure 1 shows the libraries used by the LLMs when writing code for library-agnostic tasks from BigCodeBench. All LLMs show similar preferences across the problems, with a similar distribution in the most-used libraries.
In each case, the top 3 libraries are identical (NumPy, pandas and Matplotlib) and heavily preferred, with a gap of at least 22% between themselves and the fourth most-used. It is unsurprising that these libraries are the most-preferred due to a large number of the contained problems being in the computation and visualisation domains, but they are often imported when they are not necessary. For example, the most used library, NumPy, is imported in responses to 192/301 tasks where it is not a part of the ground truth solution - showing a preference to use it over alternative options for any task.
The range of Python libraries that are available for download is expansive - in January 2025, over 7000 different open-source libraries surpassed 100,000 downloads (TopPyPIPackages2025) (this is still just a fraction of those available). However, all models use similarly low numbers of distinct libraries, ranging from 32 (Llama3.2) to 39 (Haiku3.5). Although not all of the available libraries are suitable, using less than 40 over a wide range of tasks shows a severe lack of diversity.
Case Analysis: It is clear from the results that LLMs show a preference for established libraries. To understand the extent of this bias, we perform a case analysis on libraries used in different domains, to see if there are any high-quality libraries neglected by LLMs. In particular, we use GitHub star growth over time as an objective heuristic for quality, and look for alternative libraries that show signs of higher quality but have lower usage rates.
Figure 2 shows the GitHub star growth rates for competing libraries. We observe that high-quality alternatives exist in the computation and visualisation domains. For computation problems, pandas (2010) is heavily favoured by LLMs, being used for 58% of tasks. Polars (2020) is a newer library with similar functionality (including usage of DataFrame objects) and twice as fast GitHub star growth, but is not used by LLMs at all. Similarly, Matplotlib (2011) and seaborn (2012) are heavily used by LLMs for visualisation problems, for 57% and 17% or tasks respectively. Plotly (2013) has additional functionality and signs of higher potential, but is used for only one problem and by three LLMs.
Figure 2. Case Analysis. GitHub growth statistics for libraries studied in case analysis. For each library, the accumulated GitHub stars are plotted over its lifetime (data from GitHub, December 2024). The top/bottom two graphs show case analysis for benchmark/project initialisation tasks.